how much force does it take to stretch a 10m long, 1 cm diameter steel cable by 6.0 mm?
Answer #1

We know that

Young's modulus (Y) = stress / strain

where stress = force (F) / area (A)

and

strain = change in length (L) / original length (L)

Young's modulus of steel is 2e11 N/m

so,

Y = F / A / L / L

Y = FL / AL

2e11 = F * 10 / pi * 5e-32 * 6e-3

F = 9424.77 N